# What’s the Difference? Supporting Users in Identifying the Effects of Prompt and Model Changes Through Token Patterns

### v1

@@ Abstract
Prompt engineering for large language models is challenging, as even small prompt perturbations or model changes can significantly impact the generated output texts. Existing evaluation methods, either automated metrics or human evaluation, have limitations, such as providing limited insights or being labor-intensive. We propose Spotlight, a new approach that combines both automation and human analysis. Based on data mining techniques, we automatically distinguish between random (decoding) variations and systematic differences in language model outputs. This process provides token patterns that describe the systematic differences and guide the user in manually analyzing the effects of their prompt and model changes efficiently. We create three benchmarks to quantitatively test the reliability of token pattern extraction methods and demonstrate that our approach provides new insights into established prompt data. From a human-centric perspective, through demonstration studies and a user study, we show that our token pattern approach helps users understand the systematic differences of language model outputs, and we are able to discover relevant differences caused by prompt and model changes (e.g. related to gender or culture), thus supporting the prompt engineering process and human-centric model behavior research.
@@
Michael A. Hedderich, Anyi Wang, Raoyuan Zhao, Florian Eichin, Barbara Plank LMU Munich and Munich Center for Machine Learning, Germany {mhedderich, raoyuan.zhao, feichin, b.plank}@cis.lmu.de, anyi.wang@campus.lmu.de
@@ 1 Introduction
Prompting has become a key method for controlling large language models (LLMs) Liu et al. (2023), but finding the right prompt (prompt engineering) is still challenging. While prompting allows users to influence various aspects of the generated output covering both content and style, even a semantically insignificant change in the prompt can have a substantial effect on the generated output Mizrahi et al. (2024); Shu et al. (2024). This makes finding the right prompt for a task difficult, especially in complex scenarios and when carried out by non-AI-experts Zamfirescu-Pereira et al. (2023b, a).
Different approaches exist to evaluate the generated output of LLMs and to understand the effects of prompt and model changes. Automated metrics, such as accuracy or BLEU scores Papineni et al. (2002), are fast to compute and allow for easy quantitative comparisons. However, such task-specific metrics only measure predefined criteria and just provide a simplified view on the long texts that modern LLMs generate. In contrast, human annotators can analyze texts in a nuanced and flexible way, but this process is labor-intensive Hopkins and Booth (2021) and prone to errors Wu and Aji (2023); Ghosh et al. (2024), especially in light of large amounts of LLM outputs. A core challenge when analyzing these LLM outputs is their stochastic decoding Holtzman et al. (2020). Even for the same model and prompt, two LLM outputs are likely different from each other. To distinguish systematic differences from random decoding variations, analyzing large amounts of LLM outputs is hence required when comparing two prompts or models.
In this work, we propose to combine the speed and quantification of automated methods with the flexibility of human evaluators, as visualized in Figure 1. In a nutshell, our approach works as follows: We generate a set of outputs for two different prompts or models and then automatically extract descriptions of the differences between the sets of texts. By adapting data mining techniques for this task, we leverage their reliability in identifying systematic differences. The differences are described as sets of tokens (token patterns) that offer a concise form while preserving nuances. The token patterns then guide the analysis of a human evaluator and help them understand the effects of their prompt and model changes. We call our approach Spotlight, as it shines light on Systematic Pattern differences in Output Tokens.
We see this new approach not as a replacement for existing, established evaluation techniques such as automatic metrics and full manual evaluation. Rather, we propose a middle ground and want to provide a complementary way to support users in prompt engineering and researchers in model evaluation. In contrast to existing behavioral testing like CheckList Ribeiro et al. (2020), we are interested in generated, longer LLM output texts and do not depend on a single target metric.
@@
We propose Spotlight, a token pattern based approach to identify the systematic differences caused by prompt and model changes.
@@
The code and data for this work, including the benchmarks and all token pattern results, will be made publicly available at https://github.com/mainlp/spotlight.
@@ 2.3 Human Evaluation
Human annotators can manually analyze and compare outputs, a process often considered the gold standard, but this is labor-intensive and error-prone. Experts can evaluate both the main task goals and potential, unforeseen issues, but this might require a complex evaluation process Cabrera et al. (2023b) and might be limited by resource constraints Hopkins and Booth (2021). Outsourcing the analysis to crowdworkers comes with its own challenges, such as maintaining quality Hettiachchi et al. (2022). In general, human evaluators might be inconsistent Ghosh et al. (2024), be influenced by confirmation bias Hopkins and Booth (2021) or order bias Wu and Aji (2023), or find patterns in data where none exist Ellerby and Tunney (2017). Two goals of our approach are to reduce manual effort and limit false manual conclusions by only providing token patterns to the user that indicate systematic differences.
@@ 2.4 LLM as Judge
Recently, the use of LLMs to evaluate LLM-generated outputs has gained popularity Wu et al. (2023); Akkasi et al. (2023); Gao et al. (2024); Chang et al. (2024), being automatic and often not requiring additional test data (reference-free). These approaches also suffer, however, from biases, stability, and lack of correlation with human judgments Bavaresco et al. (2024); Wang et al. (2024); Wu and Aji (2023); Deutsch et al. (2022). The risk of hallucinations Huang et al. (2025) and their low performance in math and reasoning tasks Cobbe et al. (2021); Srivatsa and Kochmar (2024) further exacerbate issues in our setting where we search for systematic differences between texts. Due to their prevalent use, we evaluate the use of LLMs as judges in our benchmarks.
@@ 2.5 Interdisciplinary Evaluations of NLP Models
Understanding the large output space of models can be seen as a sensemaking problem Pirolli and Card (2005); Gero et al. (2022), a concept from the field of human-computer interaction. User interfaces can support this process, e.g. by grouping outputs according to predefined rules Gero et al. (2022) or by highlighting significant words or sequences within each individual output text Gero et al. (2024). In these works, the user is still supposed to inspect a large set of outputs while we provide concise support to the user through succinct token patterns. Zeno Cabrera et al. (2023a) and ChainForge Arawjo et al. (2024) are platforms for visualizing and testing AI models and LLMs. Our approach could be naturally integrated into such a platform.
@@ 3 Mining Systematic LLM Output Differences
Our goal is to identify systematic differences between generated texts from two different sources, e.g., from two different prompts for the same LLM or the same prompt provided to two different LLMs. Identifying these differences is non-trivial as even two generated output texts using the same source (the same prompt and model) will likely not be identical. This is due to stochastic effects in the decoding methods which are essential for high-quality and human-like texts Holtzman et al. (2020). We therefore need to ignore the differences between the sources that are due to random decoding variations and only identify those differences that are systematic and consistent across many samples of the sources.
Identifying these differences through manual reading of large amount of textual outputs does not scale. We want to reduce the manual effort while at the same time not reducing the analysis to a simplistic metric that might miss crucial aspects. We are specifically interested in a succinct and faithful explanation of the systematic differences that can be directly linked back to and further inspected in the original data. For this, we choose sets of tokens (token patterns) as description medium. For example, a gender bias could thus be indicated by the token set (he, him) occurring much more frequently in one of the two sources. Similarly, a change in formality might be identified by token pattern phrases such as (Dear, Professor).
@@ 4 Token Pattern Discovery Methods
Spotlight is independent of a specific method to identify the token patterns and can be instantiated using different approaches. We compare counting-, data mining- and LLM-based techniques on our benchmarks. For counting and identifying systematically different token distributions, tf-idf Sparck Jones (1972) is a very established approach in NLP. We evaluate c-tf-idf by Grootendorst (2022) which is an adaptation that distinguishes between classes (here, two groups of texts).
@@ 5 Experiments with Known Ground Truth
In real use cases, the ground truth is usually not known due to the blackbox nature of LLMs, making proper evaluation difficult. Therefore, we create three benchmark datasets where the differences between the sets of texts ( and ) are injected synthetically into the texts and we thus know the exact ground truth. We make sure to obtain realistic textual settings by taking inspiration from existing NLP work.
@@ 5.1 Benchmarks
In this benchmark, we obtain by prompting an LLM for stories about doctors and filtering for only stories with male doctors. The transformation functions for this benchmark flip the gender pronouns in a percentage of the text sequences through a rule-based system. We thus ensure that only the pronouns are the systematic difference and that no other differences exist that might correlate with the gender of the doctor in the originally generated stories. For , of the stories are about male doctors, while for , we vary their amount between and to test different levels of severity of gender bias.
@@
We create by generating emails from a student to a professor. For , we use unchanged, formal texts. For , we apply a rule-based transformation function that makes the texts more informal. The set of rules includes single-token changes like adding an emoji as well as multi-token changes, e.g., replacing "Dear Professor" with a set of informal greetings like "Hi". For evaluation, we count both parts of each replacement rule as a ground truth pattern and compute precision/recall/F1 scores.
@@ 5.2 Results
The first benchmark shows that c-tf-idf, PySubgroup and the larger LLMs can identify the gender bias independent of the strength of the bias. Premise only struggles to identify the subtle bias with 60% male doctors compared to the unbiased 50%. The smaller Llama model and Cortana fail this task (Appendix Table 3).
@@
Having too many texts to analyze can also be an issue. LLMs and some data mining methods do not scale to a larger number of sequences per group, especially with LLMs quickly reaching memory or maximum token length restrictions. To raise this limit, we also experiment in Benchmark 1 with a version that is truncated to the first two sentences for each text. For the second and third benchmark, truncating is not an option, as systematic differences can occur in any part of the generated texts. The LLM methods and SPuManTE can not produce results for these benchmarks. Appendix A.1 discusses these issues in detail.
@@
The second benchmark – which contains longer patterns – is more challenging. Premise achieves an F1 score of 0.37, while the remaining methods get precision/recall/F1 scores of 0. Partial patterns are a main issue for Premise, like finding (hope, email, finds, well), missing the two pronouns (and common words) "I" and "you". The other methods fail on this task because they can not find multi-token patterns by design or do not scale to longer patterns (see Appendix Table 6 for extracted patterns). Acknowledging these limitations and evaluating with a soft score that also counts partially identified patterns gives a more nuanced picture (Appendix Figure 3, details on the metric in Appendix B.4). But even then, c-tf-idf, PySubgroup and Cortana struggle to identify these patterns. Premise still outperforms them, getting a soft precision of but missing some patterns with a soft recall of .
The third benchmark shows the influence of frequency on the methods’ performance (Appendix Figure 5). All approaches, except Cortana, are able to identify high-frequency patterns but deteriorate with decreasing pattern frequency. Only PySubgroup is still able to reliably identify medium frequency patterns.
@@ 5.3 Discussion
We next focus on evaluating the general Spotlight idea – the identification of systematic differences – as a way to support prompt engineering and uncover unexpected changes. We use Premise as pattern discovery method in the remainder of the work.
@@ 6.1 Effects of Semantically Insignificant Changes
In line with Shu et al., it is inconsistent whether the LLM output agrees or disagrees with an opinion. In our case, for variant 1 and variant 3, ca. half of the answers agree and half disagree. For variant 2, however, 92% of the answers disagree. Premise identifies corresponding patterns highlighting (False) for variant 2 when comparing it to the other two variants - for which it identifies the pattern (True) as systematically higher for them.
Analyzing the full generated texts with Spotlight uncovers further systematic differences caused by the semantically insignificant prompt changes. They have impact on the presentation, indicated by patterns such as (||, :) where 23% of outputs in variant 2 include the double bar as a structuring element (0% for the other variants). Other patterns indicate different ways the decision is motivated. Outputs of variant 3, e.g., provide a counterpoint through an (However, note) structure which occurs in 12% of the outputs (compared to only 3% and 1% for variants 1 and 2). These examples also highlight that changes can be high and low frequency. See Appendix Table 8 for details.
@@ 6.2 Effects of Personas
Shu et al. also study the effect of extending the prompt with a persona. We show that our pattern token approach identifies the same type of shifts in agreement with the previous work, but also more nuances, like a prompt-persona combination where 20% of the answers deviate from the expected binary answering style with the pattern (partially, true, some, also, may). We identify fine-grained content deviations like a "conservative" persona arguing for (traditional, roles) and (taking, responsibility) while the "open" persona’s answers mention, e.g., the (complex, multifaceted) nature of the question. Spotlight also identifies a rare but relevant failure case through the pattern (You, you, believe) where adding a specific persona can cause the LLM outputs to answer the questions for the reader instead of for "itself" as instructed in the prompt. A more detailed analysis is given in Appendix C.2.

### v2

@@ Abstract
Prompt engineering for large language models is challenging, as even small prompt perturbations or model changes can significantly impact the generated output texts. Existing evaluation methods of LLM outputs, either automated metrics or human evaluation, have limitations, such as providing limited insights or being labor-intensive. We propose Spotlight, a new approach that combines both automation and human analysis. Based on data mining techniques, we automatically distinguish between random (decoding) variations and systematic differences in language model outputs. This process provides token patterns that describe the systematic differences and guide the user in manually analyzing the effects of their prompts and changes in models efficiently. We create three benchmarks to quantitatively test the reliability of token pattern extraction methods and demonstrate that our approach provides new insights into established prompt data. From a human-centric perspective, through demonstration studies and a user study, we show that our token pattern approach helps users understand the systematic differences of language model outputs. We are further able to discover relevant differences caused by prompt and model changes (e.g. related to gender or culture), thus supporting the prompt engineering process and human-centric model behavior research.
@@
Michael A. Hedderich1, Anyi Wang1, Raoyuan Zhao1, Florian Eichin1, Jonas Fischer2, Barbara Plank1 1MaiNLP, Center for Information and Language Processing, LMU Munich Munich Center for Machine Learning (MCML) 2Department for Computer Vision and Machine Learning, Max Planck Institute for Informatics {mhedderich, raoyuan.zhao, feichin, b.plank}@cis.lmu.de, anyi.wang@campus.lmu.de,jonas.fischer@mpi-inf.mpg.de
@@ 1 Introduction
Prompting has become a key method for controlling large language models (LLMs) Liu et al. (2023), but finding the right prompt (prompt engineering) is still challenging. While prompting allows users to influence various aspects of the generated output covering both content and style, even a semantically insignificant change in the prompt can have a substantial effect on the generated output Mizrahi et al. (2024); Shu et al. (2024). This makes finding the right prompt for a task difficult, especially in complex scenarios and when carried out by non-AI-experts Zamfirescu-Pereira et al. (2023a, b).
Different approaches exist to evaluate the output generated by LLMs and to understand the effects of prompt and model changes. Automated metrics, such as accuracy or BLEU scores Papineni et al. (2002), are fast to compute and allow for easy quantitative comparisons. However, such task-specific metrics only measure predefined criteria and just provide a simplified view on the long texts that modern LLMs generate. In contrast, human annotators can analyze texts in a nuanced and flexible way, but this process is labor-intensive Hopkins and Booth (2021) and prone to errors Wu and Aji (2025); Ghosh et al. (2024), especially in light of large amounts of LLM outputs. A core challenge when analyzing LLM outputs is their stochastic decoding Holtzman et al. (2020). Even for the same model and prompt, two LLM outputs are likely different from each other. To distinguish systematic differences from random decoding variations, analyzing large amounts of LLM outputs is hence required when comparing two prompts or models.
In this work, we propose to combine the efficiency and quantification of automated methods with the flexibility of human evaluators, as visualized in Figure 1. In a nutshell, our approach works as follows: We generate a set of outputs for two different prompts or models and then automatically extract descriptions of the differences between the sets of texts. By adapting data mining techniques for this task, we leverage their reliability in identifying systematic differences. The differences are described as sets of tokens (token patterns) that offer a concise form while preserving nuances. The token patterns then guide the analysis of a human evaluator and help them understand the effects of their prompt and model changes. We call our approach Spotlight, as it shines light on Systematic Pattern differences in Output Tokens.
We see Spotlight not as a replacement for existing, established evaluation techniques such as automatic metrics and full manual evaluation. Rather, we propose a middle ground and want to provide a complementary way to support users in prompt engineering and researchers in model evaluation. In contrast to existing behavioral testing like CheckList Ribeiro et al. (2020), we are interested in generated, longer LLM output texts and do not depend on a single target metric.
@@
We propose Spotlight, a token-pattern-based approach to identify systematic differences caused by prompt and model changes.
@@
The code and data for this work, including the benchmarks and all token pattern results, are publicly available at https://github.com/mainlp/spotlight.
@@ 2.3 Human Evaluation
Human annotators can manually analyze and compare outputs, a process often considered the gold standard, but this is labor-intensive and error-prone. Experts can evaluate both the main task goals and unforeseen issues, but this might require a complex evaluation process Cabrera et al. (2023b) and might be limited by resource constraints Hopkins and Booth (2021). Outsourcing the analysis to crowdworkers comes with its own challenges, such as maintaining quality Hettiachchi et al. (2022). In general, human evaluators might be inconsistent Ghosh et al. (2024), be influenced by order bias Wu and Aji (2025), or find patterns in data where none exist Ellerby and Tunney (2017). Two goals of our approach are to reduce manual effort and limit false manual conclusions by only providing token patterns to the user that indicate systematic differences.
@@ 2.4 LLM as Judge
Recently, the use of LLMs to evaluate LLM-generated outputs has gained popularity due to them being automatic and often reference-free Wu et al. (2023); Akkasi et al. (2023); Gao et al. (2024); Chang et al. (2024); Li et al. (2024b); Xu et al. (2023). These approaches also suffer, however, from biases, instability, and lack of correlation with human judgments Bavaresco et al. (2024); Wang et al. (2024); Wu and Aji (2025); Deutsch et al. (2022); Li et al. (2024b). The risk of hallucinations Huang et al. (2025) and their low performance in math and reasoning tasks Cobbe et al. (2021); Srivatsa and Kochmar (2024) further exacerbate issues in our setting where we search for systematic differences between texts. We evaluate the suitability of LLMs as judges in our benchmarks.
@@ 2.5 Interdisciplinary Evaluations of NLP Models
Understanding the large output space of models can be seen as a sensemaking problem Pirolli and Card (2005); Gero et al. (2022), a concept from the field of human-computer interaction. User interfaces can support this process, e.g. by grouping outputs according to predefined rules Gero et al. (2022) or by highlighting significant words or sequences within each individual output text Gero et al. (2024). In these works, the user is still supposed to inspect a large set of outputs while we provide concise support to the user through succinct token patterns. Zeno Cabrera et al. (2023a) and ChainForge Arawjo et al. (2024) are platforms for visualizing and testing AI models and LLMs. Our approach could be naturally integrated into such platforms.
@@ 3 Mining Systematic LLM Output Differences
Our goal is to identify systematic differences between generated texts from two different sources, e.g., from two different prompts for the same LLM or the same prompt provided to two different LLMs. Identifying these differences is non-trivial as even two generated output texts using the same source (the same prompt and model) will likely not be identical. This is due to stochastic effects in the decoding methods which are essential for high-quality and human-like texts Holtzman et al. (2020). We, therefore, need to ignore the differences between the sources that are due to random decoding variations and only identify those differences that are systematic and consistent across many samples of the sources.
Identifying these differences through manual reading of large amounts of textual outputs does not scale. We want to reduce the manual effort while at the same time not reducing the analysis to a simplistic metric that might miss crucial aspects. Specifically, we are interested in a succinct and faithful explanation of the systematic differences that can be directly linked back to and further inspected in the original data. For this, we choose sets of tokens (token patterns) as description medium. For example, a gender bias could thus be indicated by the token set he, him occurring much more frequently in one of the two sources. Similarly, a change in formality might be identified by token pattern phrases such as Dear, Professor.
@@ 4 Token Pattern Discovery Methods
Spotlight is independent of a specific method to identify the token patterns and can be instantiated using different approaches. We compare counting-, data mining- and LLM-based techniques on our benchmarks. For counting and identifying systematically different token distributions, tf-idf Spärck Jones (1972) is a very established approach in NLP. We evaluate c-tf-idf by Grootendorst (2022) which is an adaptation that distinguishes between classes (here, two groups of texts).
@@ 5 Experiments with Known Ground Truth
In real use cases, the ground truth is usually not known due to the blackbox nature of LLMs, making proper evaluation difficult. Therefore, we create three benchmark datasets where the differences between the sets of texts ( and ) are injected synthetically into the texts, thus introducing an exact ground truth. We make sure to obtain realistic textual settings by taking inspiration from existing NLP work.
@@ 5.1 Benchmarks
In this benchmark, we obtain by prompting an LLM for stories about doctors and filtering for only stories with male doctors. The transformation functions for this benchmark flip the gender pronouns in a percentage of the text sequences through a rule-based system. We thus ensure that only the pronouns are the systematic difference and that no other differences exist that might correlate with the gender of the doctor in the originally generated stories. For , 50% of the stories are about male doctors, while for , we vary their amount between 60% and 90% to test different levels of severity of gender bias.
@@
We create by generating emails from a student to a professor. For , we use unchanged, formal texts. For , we apply a rule-based transformation function that makes the texts more informal. The set of rules includes single-token changes like adding an emoji as well as multi-token changes, e.g., replacing “Dear Professor” with a set of informal greetings like “Hi”. For evaluation, we count both parts of each replacement rule as a ground truth pattern and compute precision/recall/F1 scores.
@@ 5.2 Results
The first benchmark shows that c-tf-idf, PySubgroup, and the larger LLMs can identify the gender bias independent of the strength of the bias. Premise only fails to identify the subtle bias with 60% male doctors compared to the unbiased 50%, but identifies it for all higher bias levels. The smaller Llama model and Cortana fail this task for all bias levels.
@@
Having too many texts to analyze can also be an issue. While each text might not be that long, the sum of texts ( + ) quickly gets large. A high number of texts, however, is necessary to distinguish between random decoding variations and significant differences; including rare but significant ones like the instruction failure case we identify in Section 6.2. LLMs and some data mining methods do not scale to a larger number of sequences per group. LLMs quickly reach memory or maximum token length restrictions. Data mining methods either run into memory issues or their runtime increases to unusable levels. To raise the group size limit, we also experiment in Benchmark 1 with a version that is truncated to the first two sentences for each text (Table 1). For the second and third benchmark, truncating is not an option, as systematic differences can occur in any part of the generated texts. We exclude the LLM methods and SPuManTE for these benchmarks as they are not able an not produce results. Appendix A.1 discusses these issues in further detail.
@@
| Method | Example Patterns |
| Premise | forward, hearing, from, look, finds, well, hope, email, plz, wanna, Dear, Professor |
| c-tf-idf | to, you, and, your |
| PySubgroup | Name, !, and, ., with, !, ., I, I, and, ., I, Name, !, Thank, . |
| Cortana | Dear, Available, like, personalized, personalized, this, Sincerely, manner |
The second benchmark – which contains longer patterns – is more challenging. Premise achieves an F1 score of 0.37, while the remaining methods get precision/recall/F1 scores of 0. Partial patterns are a main issue for Premise, like finding hope, email, finds, well, missing the two pronouns (and common words) "I" and "you". The other methods fail on this task because they can not find multi-token patterns by design or do not scale to longer patterns (Table 2). Acknowledging these limitations and evaluating with a soft score that also counts partially identified patterns gives a more nuanced picture (Appendix Figure 4, details on the metric in Appendix B.4). But even then, c-tf-idf, PySubgroup and Cortana struggle to identify these patterns. Premise still outperforms them, getting a soft precision of 0.9 but missing some patterns with a soft recall of 0.7.
The third benchmark shows the influence of frequency on the methods’ performance (Figure 2). All approaches, except Cortana, are able to identify high-frequency patterns but deteriorate with decreasing pattern frequency. Only PySubgroup is still able to reliably identify medium frequency patterns.
@@ 5.3 Discussion
We next focus on evaluating the general Spotlight idea—the identification of systematic differences—as a way to support prompt engineering and uncover unexpected changes. Based on the conclusions from the benchmark, we use Premise as pattern discovery method in the remainder of the work.
@@ 6.1 Effects of Semantically Insignificant Changes
In line with Shu et al., it is inconsistent whether the LLM output agrees or disagrees with an opinion. In our case, for variant 1 and variant 3, ca. half of the answers agree and half disagree. For variant 2, however, 92% of the answers disagree. Premise identifies corresponding patterns highlighting False for variant 2 when comparing it to the other two variants - for which it identifies the pattern True as systematically higher for them.
Analyzing the full generated texts with Spotlight uncovers further systematic differences caused by the semantically insignificant prompt changes. They have impact on the presentation, indicated by patterns such as ||, : where 23% of outputs in variant 2 mirror the prompt and include the double bar as a structuring element (0% for the other variants). Other patterns indicate different ways the decision is motivated. Outputs of variant 3, e.g., provide a counterpoint through an However, note structure which occurs in 12% of the outputs (compared to only 3% and 1% for variants 1 and 2). These examples also highlight that changes can be high and low frequency. See Appendix Table 8 for details.
@@ 6.2 Effects of Personas
Shu et al. also study the effect of extending the opinion prompt questions with a persona description like “You are a conservative person, often valuing tradition, cautious about change, [...]”.
We show that our pattern token approach identifies the same type of shifts in agreement to opinion questions as the previous work, but also more nuances, like a prompt-persona combination where 20% of the answers deviate from the expected binary answering style with the pattern partially, true, some, also, may. We identify fine-grained content deviations like a "conservative" persona arguing for traditional, roles and taking, responsibility while the "open" persona’s answers mention, e.g., the complex, multifaceted nature of the question. Spotlight also identifies a rare but relevant failure case through the pattern You, you, believe where adding a specific persona can cause the LLM outputs to answer the questions for the reader instead of for "itself" as instructed in the prompt. A more detailed analysis is given in Appendix C.2.